# Entropic vector quantile regression: Duality and Gaussian case

### v1

@@ 1.1. Overview
For standard OT, entropic regularization offers various computational and analytical advantages, which has prompted extensive research activities on entropic OT [LÉO13, NUT21]. The popularity of entropic OT stems from the fact that it is amenable to efficient computation via Sinkhorn’s algorithm, for which rigorous convergence guarantees have been developed [CUT13, PC19, GN25]. Dual potentials for entropic OT, which are completely characterized by the system of functional equations, called the Schrödinger system, are smooth as long as the ground cost is smooth, which enables faster sample complexity rates than unregularized OT [CRL+20, MN19]. In addition, as the regularization parameter tends to zero, various objects of entropic OT converge to those of unregularized OT [MIK04, LÉO12, CPT23, EN24, MS25].
@@
$\inf_{\pi\in\Pi(\mu,\nu)}\big\{\mathbb{E}[\|U-\tilde{Y}\|^{2}/2]+\varepsilon\mathsf{KL}\left(\pi\,\|\,\mu\otimes\nu\right):(U,\tilde{X},\tilde{Y})\sim\pi,\,\mathbb{E}[\tilde{X}\mid U]=\mathbb{E}[X]\big\},$ (4)
where $\varepsilon>0$ is a regularization parameter and $\mathsf{KL}$ denotes the Kullback-Leibler (KL) divergence defined by
$\mathsf{KL}\left(P\,\|\,Q\right):=\begin{cases}\int\log\frac{dP}{dQ}\,dP&\text{if $P\ll Q$},\\ \infty&\text{otherwise}.\end{cases}$
@@
for suitable triplet of functions $(f,g,h)$, which directly yields strong duality and dual attainment. Our proof adapts various techniques from duality theory for entropic OT that originate from [CSI75, FG97] (see also [NUT21]). One obstacle in our proof is that (unconditional) moment constraints approximating the mean-independence constraint need not be continuous in total variation unless $X$ is compactly supported. We bypass this difficulty by imposing coercivity of the KL-divergence in the $1$-Wasserstein topology, which is shown to be satisfied under reasonably mild moment conditions.
@@ 1.4. Notation
On a Euclidean space, let $\|\cdot\|$ and $\langle\cdot,\cdot\rangle$ denote the standard Euclidean norm and inner product, respectively. For any Polish metric space $(M,d)$, we use $\mathcal{B}(M)$ to denote its Borel $\sigma$-field. Let $\mathcal{P}(M)$ denote the collection of all Borel probability measures on $M$. For any $p\in[1,\infty)$ and any fixed $x_{0}\in M$, let $\mathcal{P}_{p}(M):=\{\mu\in\mathcal{P}(M):\int d^{p}(x,x_{0})\,d\mu(x)<\infty\}$. For $\rho_{0},\rho_{1}\in\mathcal{P}_{p}(M)$, the $p$-Wasserstein distance is
@@
which defines a metric on $\mathcal{P}_{p}(M)$. For any $\mu\in\mathcal{P}(M),p\in[1,\infty]$, and $d\in\mathbb{N}$, let $L^{p}(\mu;\mathbb{R}^{d})$ denote the $L^{p}(\mu)$-space of Borel measurable mappings $M\to\mathbb{R}^{d}$. We write $L^{p}(\mu)=L^{p}(\mu;\mathbb{R})$. Finally, for $a\in\mathbb{R}$, let $a^{+}=\max\{a,0\}$ and $a^{-}=\max\{-a,0\}$.
@@ 2.1. VQR and entropic VQR
The latter condition, $\mathbb{E}[X]=0$, is for normalization and does not lose generality. Let $\mu\in\mathcal{P}(\mathbb{R}^{d_{y}})$ be a reference measure with a finite second moment.
@@
$\begin{split}\mathsf{T}^{\varepsilon}(\mu,\nu)&:=\inf_{\pi\in\Pi(\mu,\nu)}\left(\int c\,d\pi+\varepsilon\mathsf{KL}\left(\pi\,\|\,\mu\otimes\nu\right)\right)\\ &\qquad\text{subject to}\ \int_{\mathbb{R}^{d_{x}+d_{y}}}x\,d\pi_{u}(x,y)=0\ \text{$\mu$-a.e. $u$}.\end{split}$ (6)
@@
$\int c\,d\pi+\varepsilon\mathsf{KL}\left(\pi\,\|\,R\right)=\varepsilon\mathsf{KL}\left(\pi\,\|\,\tilde{R}\right)-\varepsilon\log\alpha.$
@@ 2.2. Duality
$\iota^{\varepsilon}(f,g,h):=\varepsilon\int_{\mathbb{R}^{d_{y}}\times\mathbb{R}^{d_{x}+d_{y}}}e^{\frac{1}{\varepsilon}\left(f(u)+\langle g(u),x\rangle+h(x,y)-c(u,y)\right)}\,dR(u,x,y)-\varepsilon$
@@
We first establish weak duality.
@@
For strong duality and dual attainment, we will make the following additional assumption. Recall that $\tilde{R}$ is a probability measure on $\mathbb{R}^{d_{y}}\times\mathbb{R}^{d_{x}+d_{y}}$ defined in (8).
@@
A $\beta$-sub-Weibull variable $\xi$ is often called sub-Gaussian if $\beta=2$ and sub-exponential if $\beta=1$. We say that a random vector $Z$ or its law is $\beta$-sub-Weibull if $\|\|Z\|\|_{\psi_{\beta}}<\infty$. The next lemma provides simple sufficient conditions for Assumption 2.1 (ii). The proof relies on a weighted Pinsker inequality from [BV05].
Lemma 2.1.
Assumption 2.1 (ii) holds if either one of the following conditions is satisfied for some $\beta>1$: (i) $\mu$ and $\nu$ are both $\beta$-sub-Weibull; or (ii) $\mu$ is compactly supported and $X$ is $\beta$-sub-Weibull.
@@
For given $(f,g,h)\in L^{1}(\mu)\times L^{1}(\mu;\mathbb{R}^{d_{x}})\times L^{1}(\nu)$, they solve the dual problem (9) if and only if they satisfy
@@ 2.3. Regularity of dual potentials
In this section, we assume the existence of dual potentials and study their regularity via the Schrödinger-like system (12)–(14). Throughout this section, we choose versions of dual potentials $(f^{\varepsilon},g^{\varepsilon},h^{\varepsilon})$ so that (12) and (14) hold for all $u\in\mathbb{R}^{d_{y}}$ and $(x,y)\in\mathbb{R}^{d_{x}+d_{y}}$, respectively; see Remark 2.1. Furthermore, let $\mathcal{U}\subset\mathbb{R}^{d_{y}}$ denote the support of $\mu$, and let $\mathcal{X}\subset\mathbb{R}^{d_{x}},\mathcal{Y}\subset\mathbb{R}^{d_{y}}$ denote the supports of $X,Y$, respectively. We will assume throughout this section that $\mathcal{U}$ is compact.
@@ 4. Proofs for Section 2
Uniqueness follows from strict convexity of $\mathsf{KL}\left(\cdot\,\|\,\mu\otimes\nu\right)$, so we prove the existence. We first rewrite the primal problem (6) as a weak OT problem [BPR+25]. Since $\mathsf{KL}\left(\pi\,\|\,\mu\otimes\nu\right)=\int_{\mathbb{R}^{d_{y}}}\mathsf{KL}\left(\pi_{u}\,\|\,\nu\right)\,d\mu(u)$ [PW25, Theorem 2.15], we have
$\int c\,d\pi+\varepsilon\mathsf{KL}\left(\pi\,\|\,\mu\otimes\nu\right)=\int_{\mathbb{R}^{d_{y}}}\left(\int_{\mathbb{R}^{d_{x}+d_{y}}}c(u,y)\,d\pi_{u}(x,y)+\varepsilon\mathsf{KL}\left(\pi_{u}\,\|\,\nu\right)\right)\,d\mu(u).$
@@
$C(u,\rho):=\int_{\mathbb{R}^{d_{x}+d_{y}}}c(u,y)\,d\rho(x,y)+\varepsilon\mathsf{KL}\left(\rho\,\|\,\nu\right)+\chi_{\mathcal{C}}(\rho).$ (18)
We will apply Theorem 2.2 (i) in [BPR+25] to establish the desired claim. Endow $\mathcal{P}_{2}(\mathbb{R}^{d_{x}+d_{y}})$ with the $2$-Wasserstein distance. We need to verify that the cost $C$ is Borel, and that $\rho\mapsto C(u,\rho)$ is convex and lower semicontinuous. The mapping $(u,\rho)\mapsto\int_{\mathbb{R}^{d_{x}+d_{y}}}c(u,y)\,d\rho(x,y)$ is jointly continuous and linear in $\rho$, and $\rho\mapsto\mathsf{KL}\left(\rho\,\|\,\nu\right)$ is convex and lower semicontinuous by the Donsker-Varadhan variational representation (cf. Theorem 4.6 in [PW25]). The set $\mathcal{C}$ is convex and closed in $\mathcal{P}_{2}(\mathbb{R}^{d_{x}+d_{y}})$, so $\chi_{\mathcal{C}}$ is convex and lower semicontinuous. As such, applying Theorem 2.2 (i) in [BPR+25], we obtain the desired claim. ∎
@@
Pick any feasible coupling $\pi\in\mathcal{Q}$ with $\mathsf{KL}\left(\pi\,\|\,R\right)<\infty$, and let $p_{\pi}:=\frac{d\pi}{dR}$. Then $\int p_{\pi}\,dR=\int d\pi=1$. Let $F(u,x,y):=f(u)+\langle g(u),x\rangle+h(x,y)$. Observe that
@@
We shall verify that $\langle g(u),x\rangle\in L^{1}(\pi)$. Since the left-hand side and the second term on the right-hand side on (19) are integrable under $R$, we have $(F-c)^{+}p_{\pi}\in L^{1}(R)$, i.e., $(F-c)^{+}\in L^{1}(\pi)$. Observe that
@@
The right-hand side is integrable under $\pi$, so that $(\langle g(u),x\rangle)^{+}\in L^{1}(\pi)$. Now, since $\pi$ is a coupling for $(\mu,\nu)$, $\|x\|$ is integrable under $\pi$, so that by Fubini’s theorem,
$\int_{\mathbb{R}^{d_{x}+d_{y}}}\|x\|\,d\pi_{u}(x,y)<\infty\ \text{$\mu$-a.e. $u$}$
@@
$\varepsilon\mathsf{KL}\left(\pi\,\|\,R\right)-\varepsilon\geq\left(\int f\,d\mu+\int h\,d\nu-\int c\,d\pi\right)-\iota^{\varepsilon}(f,g,h)-\varepsilon.$
@@
The proof relies on the following lemma.
@@
Let $(M,\|\cdot\|)$ be a finite-dimensional normed space, and let $Q\in\mathcal{P}(M)$ be $\beta$-sub-Weibull for some $\beta>1$. Then the mapping $P\mapsto\mathsf{KL}\left(P\,\|\,Q\right)$ is coercive in $\mathsf{W}_{1}$.
@@
Since $Q$ is $\beta$-sub-Weibull for some $\beta>1$, Corollary 2.3 in [BV05] yields that there exists a finite constant $K$ such that
$\mathsf{W}_{1}(P,Q)\leq\mathsf{W}_{\beta}(P,Q)\leq K\left(\mathsf{KL}\left(P\,\|\,Q\right)^{\frac{1}{\beta}}+\mathsf{KL}\left(P\,\|\,Q\right)^{\frac{1}{2\beta}}\right),\ \forall P\in\mathcal{P}(M).$ (20)
Fix $0<a<\infty$ and consider the sublevel set
$\mathcal{L}_{a}=\{P\in\mathcal{P}(M):\mathsf{KL}\left(P\,\|\,Q\right)\leq a\}.$
Pick any sequence $P_{n}\in\mathcal{L}_{a}$. By inequality (20), $\mathsf{W}_{1}(P_{n},Q)$ is bounded, so that $\int\|x\|dP_{n}(x)$ is bounded in $n$. This implies that $P_{n}$ is uniformly tight and $P_{n}\to P$ weakly along a subsequence for some $P\in\mathcal{P}(M)$. We will verify that $P_{n}$ is uniformly integrable, which implies that $P_{n}\to P$ in $\mathsf{W}_{1}$ and
$\mathsf{KL}\left(P\,\|\,Q\right)\leq\liminf_{n}\mathsf{KL}\left(P_{n}\,\|\,Q\right)\leq a$
by lower semicontinuity of $\mathsf{KL}\left(\cdot\,\|\,Q\right)$ in $\mathsf{W}_{1}$.
Establishing uniform integrability of $P_{n}$ requires some effort. Pick any $P\in\mathcal{L}_{a}$. By the data processing inequality (cf. Theorem 7.4 in [PW25]), one has
$\mathsf{KL}\left(P\,\|\,Q\right)\geq\mathsf{kl}(P(A)||Q(A)),\ A\subset M,$
@@
$\mathsf{kl}(p||q)=p\log\frac{p}{q}+(1-p)\log\frac{1-p}{1-q}$
@@
$\mathsf{kl}(p||q)=p\log\frac{1}{q}+\underbrace{\big[p\log p+(1-p)\log(1-p)\big]}_{\geq-\log 2}+\underbrace{(1-p)\log\frac{1}{1-q}}_{\geq 0}\geq p\log\frac{1}{q}-\log 2.$

### v2

@@ 1.1. Overview
For standard OT, entropic regularization offers various computational and analytical advantages, which has prompted extensive research activities on entropic OT [LÉO13, NUT21]. The popularity of entropic OT stems from the fact that it is amenable to efficient computation via a dual block coordinate ascent algorithm, called Sinkhorn’s algorithm, for which rigorous convergence guarantees have been developed [SIN67, FL89, CUT13, PC19, GN25]. Dual potentials for entropic OT, which are completely characterized by the system of functional equations, called the Schrödinger system, are smooth provided that the ground cost is smooth and marginals have sufficiently light tails, which enables faster sample complexity rates than unregularized OT [GCB+19, MN19, dGL+23, GLN22, GKR+24]. In addition, as the regularization parameter tends to zero, various objects of entropic OT converge to those of unregularized OT [MIK04, LÉO12, NW22, CPT23, EN24, MS25]
@@
$\inf_{\pi\in\Pi(\mu,\nu)}\big\{\mathbb{E}[\|U-\tilde{Y}\|^{2}/2]+\varepsilon\mathsf{KL}(\pi\,\|\,\mu\otimes\nu):(U,\tilde{X},\tilde{Y})\sim\pi,\,\mathbb{E}[\tilde{X}\mid U]=\mathbb{E}[X]\big\},$ (4)
where $\varepsilon>0$ is a regularization parameter and $\mathsf{KL}$ denotes the Kullback-Leibler (KL) divergence (or relative entropy) defined by
$\mathsf{KL}(P\,\|\,Q):=\begin{cases}\int\log\frac{dP}{dQ}\,dP&\text{if $P\ll Q$},\\ \infty&\text{otherwise}.\end{cases}$
@@
for suitable triplet of functions $(f,g,h)$, which yields strong duality and dual attainment. Our proof adapts various techniques from duality theory for entropic OT that originate from [CSI75, FÖL88, FG97] (see also [NUT21]). One obstacle in our proof is that (unconditional) moment constraints approximating the mean-independence constraint need not be continuous in total variation unless $X$ is compactly supported. We bypass this difficulty by imposing coercivity of the KL-divergence in the $1$-Wasserstein topology with respect to a modified Euclidean metric, which appears to be new and can be adapted to handle different conditional moment constraints. In addition, we show that the preceding coercivity assumption holds under a reasonably mild moment condition on $X$. As another difficulty compared with standard entropic OT, one function $g$ appears in the dual problem through the interaction term, $\langle g(u),x\rangle$, between $u$ (the “input” variable) and $x$ (part of the “output” variables). As such, a more careful measure-theoretic argument is needed to separately construct dual potentials in a measurable way; see the discussion above Lemma 4.4.
@@ 1.4. Notation
On a Euclidean space, let $\|\cdot\|$ and $\langle\cdot,\cdot\rangle$ denote the standard Euclidean norm and inner product, respectively. For any Polish metric space $(M,d)$, we use $\mathcal{B}(M)$ to denote its Borel $\sigma$-field. Let $\mathcal{P}(M)$ denote the collection of all Borel probability measures on $M$. When $M$ is a finite-dimensional Euclidean space, we consider the standard Euclidean metric, unless otherwise stated. For any $p\in[1,\infty)$ and any fixed $x_{0}\in M$, let $\mathcal{P}_{p}(M):=\{\mu\in\mathcal{P}(M):\int d^{p}(x,x_{0})\,d\mu(x)<\infty\}$. For $\rho_{0},\rho_{1}\in\mathcal{P}_{p}(M)$, the $p$-Wasserstein distance is
@@
which defines a metric on $\mathcal{P}_{p}(M)$. For any $\mu\in\mathcal{P}(M),p\in[1,\infty]$, and $d\in\mathbb{N}$, let $L^{p}(\mu;\mathbb{R}^{d})$ denote the $L^{p}(\mu)$-space of Borel measurable mappings $M\to\mathbb{R}^{d}$. We write $L^{p}(\mu)=L^{p}(\mu;\mathbb{R})$. Finally, for $a\in\mathbb{R}$, let $a^{+}=\max\{a,0\}$ and $a^{-}=\max\{-a,0\}$; for $a,b\in\mathbb{R}$, we use the notation $a\wedge b=\min\{a,b\}$.
@@ 2.1. VQR and entropic VQR
The latter condition, $\mathbb{E}[X]=0$, is for normalization and does not lose generality. Let $\mu\in\mathcal{P}(\mathbb{R}^{d_{y}})$ be a reference measure with a finite second moment. We do not assume that $\mu$ is absolutely continuous.
@@
$\int\varphi\,d\pi=\int_{\mathbb{R}^{d_{y}}}\left(\int_{\mathbb{R}^{d_{x}+d_{y}}}\varphi(u,x,y)\,d\pi_{u}(x,y)\right)\,d\mu(u).$ (5)
See Chapter 10.2 in [DUD02] for (regular) conditional distributions. With this notation, the VQR problem (3) reads as
@@
$\begin{split}\mathsf{T}^{\varepsilon}(\mu,\nu)&:=\inf_{\pi\in\Pi(\mu,\nu)}\left(\int c\,d\pi+\varepsilon\mathsf{KL}(\pi\,\|\,\mu\otimes\nu)\right)\\ &\qquad\text{subject to}\ \int_{\mathbb{R}^{d_{x}+d_{y}}}x\,d\pi_{u}(x,y)=0\ \text{$\mu$-a.e. $u$}.\end{split}$ (7)
@@
$\int c\,d\pi+\varepsilon\mathsf{KL}(\pi\,\|\,R)=\varepsilon\mathsf{KL}(\pi\,\|\,\tilde{R})-\varepsilon\log\alpha.$
@@ 2.2. Duality
$\iota^{\varepsilon}(f,g,h):=\varepsilon\int_{\mathbb{R}^{d_{y}}\times\mathbb{R}^{d_{x}+d_{y}}}e^{\frac{1}{\varepsilon}\left(f(u)+\langle g(u),x\rangle+h(x,y)-c(u,y)\right)}\,dR(u,x,y)-\varepsilon.$
@@
We first establish weak duality. The proof is standard except for one place where we verify that $\langle g(u),x\rangle\in L^{1}(\pi)$ for any $\pi\in\mathcal{Q}$ with $\mathsf{KL}(\pi\,\|\,R)<\infty$, provided that $\iota^{\varepsilon}(f,g,h)<\infty$.
@@
For strong duality and dual attainment, we will make the following additional assumption. Recall that $\tilde{R}$ is a probability measure on $\mathbb{R}^{d_{y}}\times\mathbb{R}^{d_{x}+d_{y}}$ defined in (9). We consider the $1$-Wasserstein distance for a modified metric on $\mathbb{R}^{d_{y}}\times\mathbb{R}^{d_{x}+d_{y}}$ different from the Euclidean one. Define
$\tilde{\mathsf{d}}\big((u,x,y),(u^{\prime},x^{\prime},y^{\prime})\big):=\|u-u^{\prime}\|\wedge 1+\|x-x^{\prime}\|+\|y-y^{\prime}\|\wedge 1$
for $(u,x,y),(u^{\prime},x^{\prime},y^{\prime})\in\mathbb{R}^{d_{y}}\times\mathbb{R}^{d_{x}+d_{y}}$. The metric $\tilde{\mathsf{d}}$ induces the same topology as the Euclidean one. We denote by $\tilde{\mathsf{W}}_{1}$ the corresponding $1$-Wasserstein distance,
$\tilde{\mathsf{W}}_{1}(P,Q):=\inf_{\gamma\in\Pi(P,Q)}\int\tilde{\mathsf{d}}\,d\gamma,$
which defines a metric on
$\tilde{\mathcal{P}}_{1}:=\left\{P\in\mathcal{P}(\mathbb{R}^{d_{y}}\times\mathbb{R}^{d_{x}+d_{y}}):\int\tilde{\mathsf{d}}(\cdot,0)\,dP<\infty\right\}.$ (11)
By definition, $P\in\tilde{\mathcal{P}}_{1}$ if and only if $\mathbb{E}[\|\tilde{X}\|$$\infty$ for $(U,\tilde{X},\tilde{Y})\sim P$. Observe that $P_{n}\to P$ in $\tilde{\mathsf{W}}_{1}$ if and only if, for any continuous function $\varphi:\mathbb{R}^{d_{y}}\times\mathbb{R}^{d_{x}+d_{y}}\to\mathbb{R}$ with $|\varphi(u,x,y)|\leq K(1+\|x\|)$ for $(u,x,y)\in\mathbb{R}^{d_{y}}\times\mathbb{R}^{d_{x}+d_{y}}$ for some finite constant $K$, it holds that $\int\varphi\,dP_{n}\to\int\varphi dP$. See Theorem 6.9 in [VIL09]. In particular, the topology induced by $\tilde{\mathsf{W}}_{1}$ is stronger than the weak topology.
@@
$\big\{Q\in\tilde{\mathcal{P}}_{1}:\mathsf{KL}(Q\,\|\,\tilde{R})\leq a\big\}$
is compact for the $\tilde{\mathsf{W}}_{1}$-topology.
@@
Lemma 2.1.
Assumption 2.1 (ii) holds if
$\mathbb{E}\big[e^{\alpha^{\prime}\|X\|}\big]<\infty,\ \forall\alpha^{\prime}>0.$ (12)
Remark 2.1.
A few remarks are in order.
  1. (i)
    Recall that a real-valued random variable $\xi$ or its law is called $\beta$-sub-Weibull for some $\beta\in(0,\infty)$ if
@@
    We say that a random vector $Z$ or its law is $\beta$-sub-Weibull if $\|\|Z\|\|_{\psi_{\beta}}<\infty$. Condition (12) is satisfied if $X$ is $\beta$-sub-Weibull for some $\beta>1$.
  2. (ii)
    Condition (12) is not necessary for Assumption 2.1 (ii) to hold. See Remark 4.2 below.
@@
Suppose that Assumption 2.1 holds. For given $(f,g,h)\in L^{1}(\mu)\times L^{1}(\mu;\mathbb{R}^{d_{x}})\times L^{1}(\nu)$, they solve the dual problem (10) if and only if they satisfy
@@ 2.3. Regularity of dual potentials
In this section, we assume the existence of dual potentials satisfying the Schrödinger-like system (15)–(17) and study their regularity. Throughout this section, we choose versions of dual potentials $(f^{\varepsilon},g^{\varepsilon},h^{\varepsilon})$ so that (15) and (17) hold for all $u\in\mathbb{R}^{d_{y}}$ and $(x,y)\in\mathbb{R}^{d_{x}+d_{y}}$, respectively; see Remark 2.2. Furthermore, let $\mathcal{U}\subset\mathbb{R}^{d_{y}}$ denote the support of $\mu$, and let $\mathcal{X}\subset\mathbb{R}^{d_{x}},\mathcal{Y}\subset\mathbb{R}^{d_{y}}$ denote the supports of $X,Y$, respectively. We will assume throughout this section that $\mathcal{U}$ is compact.
@@ 4. Proofs for Section 2
We recall some properties of the KL divergence. Let $M$ be a Polish metric space. For $P,Q\in\mathcal{P}(M)$, $\mathsf{KL}(P\,\|\,Q)\in[0,\infty]$ and $\mathsf{KL}(P\,\|\,Q)=0$ if and only if $P=Q$. In addition, the mapping $P\mapsto\mathsf{KL}(P\,\|\,Q)$ is convex, lower semicontinuous for the weak topology (and any stronger topology), and strictly convex on its domain. The lower semicontinuity follows by the Donsker-Varadhan variational representation (cf. Theorem 4.6 in [PW25]).
@@
Uniqueness follows from strict convexity of $\mathsf{KL}(\cdot\,\|\,\mu\otimes\nu)$ on its domain, so we prove the existence. In fact, it is not very hard to verify that $\mathcal{Q}$ is weakly compact (as each $\pi\in\mathcal{Q}$ has fixed marginals and $X$ has finite expectation), which yields the existence.
We provide an alternative argument by rewriting the primal problem (7) as a weak OT problem [BPR+25]. Since $\mathsf{KL}(\pi\,\|\,\mu\otimes\nu)=\int_{\mathbb{R}^{d_{y}}}\mathsf{KL}(\pi_{u}\,\|\,\nu)\,d\mu(u)$ [PW25, Theorem 2.15], we have
$\int c\,d\pi+\varepsilon\mathsf{KL}(\pi\,\|\,\mu\otimes\nu)=\int_{\mathbb{R}^{d_{y}}}\left(\int_{\mathbb{R}^{d_{x}+d_{y}}}c(u,y)\,d\pi_{u}(x,y)+\varepsilon\mathsf{KL}(\pi_{u}\,\|\,\nu)\right)\,d\mu(u).$
@@
$C(u,\rho):=\int_{\mathbb{R}^{d_{x}+d_{y}}}c(u,y)\,d\rho(x,y)+\varepsilon\mathsf{KL}(\rho\,\|\,\nu)+\chi_{\mathcal{C}}(\rho).$ (21)
We will apply Theorem 2.2 (i) in [BPR+25] to establish the desired claim. Endow $\mathcal{P}_{2}(\mathbb{R}^{d_{x}+d_{y}})$ with the $2$-Wasserstein distance. We need to verify that the cost $C$ is Borel, and that $\rho\mapsto C(u,\rho)$ is convex and lower semicontinuous. The mapping $(u,\rho)\mapsto\int_{\mathbb{R}^{d_{x}+d_{y}}}c(u,y)\,d\rho(x,y)$ is jointly continuous and linear in $\rho$, and $\rho\mapsto\mathsf{KL}(\rho\,\|\,\nu)$ is convex and lower semicontinuous. The set $\mathcal{C}$ is convex and closed in $\mathcal{P}_{2}(\mathbb{R}^{d_{x}+d_{y}})$, so $\chi_{\mathcal{C}}$ is convex and lower semicontinuous. As such, applying Theorem 2.2 (i) in [BPR+25], we obtain the desired claim. ∎
@@
Pick any feasible coupling $\pi\in\mathcal{Q}$ with $\mathsf{KL}(\pi\,\|\,R)<\infty$, and let $p_{\pi}:=\frac{d\pi}{dR}$. Then $\int p_{\pi}\,dR=\int d\pi=1$. Let $F(u,x,y):=f(u)+\langle g(u),x\rangle+h(x,y)$. Observe that
@@
We shall verify that $\langle g(u),x\rangle\in L^{1}(\pi)$. Since the left-hand side and the second term on the right-hand side of (22) are integrable under $R$, we have $(F-c)^{+}p_{\pi}\in L^{1}(R)$, i.e., $(F-c)^{+}\in L^{1}(\pi)$. Observe that
@@
The right-hand side is integrable under $\pi$, so that $(\langle g(u),x\rangle)^{+}\in L^{1}(\pi)$. Now, since $\pi$ is a coupling for $(\mu,\nu)$, $\|x\|$ is integrable under $\pi$, so that (cf. (5))
$\int_{\mathbb{R}^{d_{x}+d_{y}}}\|x\|\,d\pi_{u}(x,y)<\infty\ \text{$\mu$-a.e. $u$.}$
@@
$\varepsilon\mathsf{KL}(\pi\,\|\,R)-\varepsilon\geq\left(\int f\,d\mu+\int h\,d\nu-\int c\,d\pi\right)-\iota^{\varepsilon}(f,g,h)-\varepsilon.$
@@
Lemma 2.1 directly follows from the next lemma and the fact that $e^{-c/\varepsilon}$ is bounded.
@@
Let $(M,d)$ be a Polish metric space, and let $Q\in\mathcal{P}(M)$ be satisfy
$\int e^{\alpha d(x,x_{0})}\,dQ(x)<\infty,\ \forall\alpha>0,$ (23)
for some $x_{0}\in M$. Then the mapping $P\mapsto\mathsf{KL}(P\,\|\,Q)$ is coercive in $\mathsf{W}_{1}$.
There is a partial converse to Lemma 4.1; see Appendix A.
@@
The lemma (indirectly) follows from Theorem 2.1 in [WWW10] that establishes the $\mathsf{W}_{p}$-version of Sanov’s theorem on large deviations for any $p\in[1,\infty)$. We provide a more direct proof using the weighted Pinsker inequality of Bolley and Villani [BV05], which may be of independent interest.
Fix any $0<a<\infty$ and consider the sublevel set
$\mathcal{L}_{a}=\{P\in\mathcal{P}_{1}(M):\mathsf{KL}(P\,\|\,Q)\leq a\}.$
We first show that $\mathcal{L}_{a}$ is weakly precompact. [^2] Weak precompactness of $\mathcal{L}_{a}$ also follows from the (classical) Sanov theorem (cf. Theorem 3.2.17 in [DS01]). Pick any $P\in\mathcal{L}_{a}$. By the data processing inequality (cf. Theorem 7.4 in [PW25]), one has
$\mathsf{KL}(P\,\|\,Q)\geq\mathsf{kl}(P(A)\,\|\,Q(A)),\ A\subset M,$
@@
$\mathsf{kl}(p\,\|\,q)=p\log\frac{p}{q}+(1-p)\log\frac{1-p}{1-q}$
@@
$\mathsf{kl}(p\,\|\,q)=p\log\frac{1}{q}+\underbrace{\big[p\log p+(1-p)\log(1-p)\big]}_{\geq-\log 2}+\underbrace{(1-p)\log\frac{1}{1-q}}_{\geq 0}\geq p\log\frac{1}{q}-\log 2.$